OpenStudy (anonymous):

Find the next term in the sequence 128,64,32....

OpenStudy (ivettef365):

Hint: divide 128/64

OpenStudy (anonymous):

164/64=2 64/2=32 32/2=16

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

thanks

OpenStudy (ivettef365):

yw :)